QUESTION 1 If the number of parameters in a problem is m and the number of primary dimensions is 3, then the number of resultFluid Mechanics Chapter 7 Intro
Answer #1

Question 18 Il number of paremete es in the ism & the number of primary dimension is 3 , then the resultant pi terms is m-3 .ath=o o a=-b - a+c= 0 = a c o Let al = n where foonER ab=-2 for n=1, IT = VET Question 3 - Tw=fC v Rh, mal, D, e) Torell No.